# NocPlace: Nocturnal Visual Place Recognition Using Generative and Inherited Knowledge Transfer

### 27 Feb 2024 v1

@@ Abstract
Visual Place Recognition (VPR) is crucial in computer vision, aiming to retrieve database images similar to a query image from an extensive collection of known images. However, like many vision-related tasks, learning-based VPR often experiences a decline in performance during nighttime due to the scarcity of nighttime images. Specifically, VPR needs to address the cross-domain problem of night-to-day rather than just the issue of a single nighttime domain. In response to these issues, we present NocPlace, which leverages a generated large-scale, multi-view, nighttime VPR dataset to embed resilience against dazzling lights and extreme darkness in the learned global descriptor. Firstly, we establish a day-night urban scene dataset called NightCities, capturing diverse nighttime scenarios and lighting variations across 60 cities globally. Following this, an unpaired image-to-image translation network is trained on this dataset. Using this trained translation network, we process an existing VPR dataset, thereby obtaining its nighttime version. The NocPlace is then fine-tuned using night-style images, the original labels, and descriptors inherited from the Daytime VPR model. Comprehensive experiments on various nighttime VPR test sets reveal that NocPlace considerably surpasses previous state-of-the-art methods.
@@ 1 Introduction
Figure 1: A demo for NocPlace. The center of this figure presents a map of San Francisco. 2.8 million daytime database images and 466 nighttime query images from the SF-XL dataset are projected onto this map via GPS tags, corresponding to the dense cyan-blue and sparse red dots, respectively. Despite the dazzling lights and extreme darkness, our method achieves higher recall@N than previous methods.
Visual Place Recognition (VPR) is a foundational task in computer vision [5, 24, 45, 52, 51, 65] and intelligent robotics [20, 32, 34, 36, 39, 58, 33, 61]. The goal of VPR is to retrieve similar database images to a given query image by comparing against a vast set of known database images. Existing literature identifies several challenges for VPR, including database scale [9], viewpoint shifts [8, 11], repeated structures [51], structural modifications [23], occlusions [5], visual scale differences [17], illumination changes [4, 12, 34, 32, 44, 50, 52, 58, 65], and seasonal transitions [36, 49]. Recent VPR approaches primarily leverage large-scale datasets [9, 35, 37, 52, 60, 3], and utilize neural networks to map images into an embedding space, thereby effectively distinguishing images from different locations.
However, night-to-day VPR is constrained by the current training datasets. On the one hand, recent studies [5, 9, 3] leveraging the Google Street View (GSV) download interface have produced large-scale multi-view VPR datasets covering nearly all challenges except those in nighttime conditions. On the other hand, while some datasets specifically targeting night-to-day VPR datasets [34, 36] have been proposed and shown the benefit of incorporating nighttime data for cross-domain recognition, their scale or data acquisition methods impede the training of a universally night-to-day VPR model. It is worth acknowledging that constructing a nighttime urban street-scape dataset aligned with commercial offerings like GSV is challenging.
Given the lack of comprehensive end-to-end night-to-day VPR solutions, image enhancement algorithms (IE) [69, 67, 22, 21] and night-to-day Image-to-Image (I2I) translation [4] may be a potential approach. These methods, however, present several limitations: i) Introducing an auxiliary model significantly increases computational demands. ii) Most low-light enhancement models rely on images with diverse exposure rather than authentic night-to-day transitions, and much of the data is not even sourced outdoors. iii) Imperfect Night-to-Day translations might adversely affect VPR performance [4].
In this paper, we reverse the night-to-day thinking and study how to transfer night information into neural networks directly. Further, we study the use of daytime knowledge to guide the training of nighttime models. Our method can improve the performance of nighttime visual place recognition without consuming additional computing resources.
@@
  • We propose NocPlace, a scalable methodology for Nocturnal Visual Place Recognition, which significantly narrows the VPR performance gap between daytime and nighttime domains.
  • We propose NightCities, an unpaired day-night urban scene image dataset. We utilize this dataset and an I2I translation network to transfer night knowledge to an existing large-scale multi-view daytime VPR dataset.
  • We propose the inherited knowledge transfer which employs descriptors inherited from previous VPR models to guide the training of new VPR models.
  • Extensive experiments demonstrate that NocPlace substantially outperforms previous methods. Specifically, for certain datasets, our VGG16-based model excels over state-of-the-art ResNet50-based models.
  • The source code and trained models for NocPlace are available at TBD.
@@ 2 Related Work
Visual Place Recognition can be categorized along three pivotal axes: model architecture, loss function, and dataset. With regard to model architecture, Convolutional Neural Networks (CNNs) [5] and Visual Transformers (ViTs) [53, 56, 10] emerge as the primary choices for feature extraction. These are often followed by pooling or aggregation mechanisms like NetVLAD [5, 41]. In the realm of loss functions, triplet and classification losses are frequently employed to optimize Euclidean margins for enhanced feature embeddings [10, 5]. Ge et al. utilized triplet loss and network output scores as self-supervised indicators, attaining state-of-the-art outcomes via iterative training. However, triplet loss encounters challenges with hard negative sample mining. In response, Berton et al. [9, 11] incorporated Large Margin Cosine (LMC) Loss [55] in VPR, showcasing its superiority over triplet loss. On the dataset front, we observe two primary classifications: self-curated datasets and web-sourced ones. The former, such as [20, 34, 36], often exhibit more minor scales, fixed frontal camera view, and little nighttime data. The latter, however, can be sub-categorized into datasets like Google Landmark [41], derived from Google Images, and those like SF-XL [9] and Pittsburgh [5], sourced from Google Street View [5, 1, 3, 9]. It is worth noting that the latter dataset contains almost no night data. Therefore, it can be inferred that the performance of existing VPR models in nighttime scenarios benefits from data augmentation [10] during the training phase and the generalization capability during the inference phase.
@@
Night Computer Vision encompasses traditional downstream tasks using nighttime imagery, which are primarily divided into two-stage and one-stage approaches. The two-stage approaches [4, 62] first transform nighttime images into their daytime counterparts before executing the downstream tasks. In contrast, one-stage methods directly utilize raw nighttime data for training [14, 65]. For example, Anoosheh et al. [4] introduced a CycleGAN-based method to transform nighttime images for retrieval-based localization. Cui et al. [14] developed a method for reverse ISP and dark processing, leading to an end-to-end model for detecting objects in dark settings. Additionally, Xu et al. [62] presented a GAN-centric strategy for semantic segmentation in autonomous driving scenarios using transformed nighttime imagery. However, within the VPR domain, the two-stage methods are criticized for their limited generalizability [4] and heightened computational demands. Meanwhile, the one-stage approach faces challenges due to the lack of extensive training datasets. In response, we propose a pioneering framework by converting daytime images to nighttime versions for enhanced VPR training.
@@
Image-to-Image Translation pertains to the intersection of vision and graphics, aiming to establish a mapping between an input and an output image through a designated training set. These methods are primarily categorized based on the nature of their training sets: pixel-level matched I2I [27] or unpaired I2I [71, 64]. Securing extensive pixel-level correspondent day-night image pairs in real-world scenarios is notably challenging. Hence, our focus is on the latter training set type. Two predominant strategies for unpaired image-to-image I2I translation exist: bi-directional mapping and uni-directional mapping. Bi-directional mapping, encompassing CycleGAN [71] and DualGAN [64], is anchored on the cycle-consistency constraint, ensuring the transposed image can undergo reverse mapping for reconstruction. The bi-directional mapping can be overly restrictive, prompting the evolution of uni-directional mapping techniques. Some techniques employ geometry distance for content retention. For instance, DistanceGAN [16] conserves inter-image distances within domains, while GCGAN [7] upholds geometry consistency between inputs and outputs. The CUT method [38] amplifies mutual information between input and output via contrastive learning. Recently, the diffusion model [70] has showcased exceptional efficacy in addressing I2I translation challenges, albeit with significant computational demands. After an exhaustive evaluation of various strategies, our day-to-night I2I translation model was developed, rooted in the NEGCUT framework [57].
@@
Knowledge Transfer can be categorized in various ways based on the problems it addresses, such as domain adaptation [66], multi-task learning [68], meta-learning [54], and more. This paper focuses on domain adaptation, aiming to adapt models trained in one domain (source) to perform optimally in another domain (target) where the data distribution might differ. Previous Night-to-Day VPR studies primarily focused on labeled target domains, a relatively straightforward problem [61, 63, 40]. Cao et al. [40] incorporated Canny edge descriptors and the Wasserstein distance measure for domain adaptation in VPR tasks. When dealing with labeled source and unlabeled target domains, prior research can be broadly divided into two categories: i) generation of target domain data, as exemplified by SimGAN [47], which transfers samples from the source to the target domain while preserving consistent labels. ii) adversarial learning of domain-invariant features. Ganin et al. introduced DANN [18], which learns domain-invariant feature distributions by minimizing the source domain classification loss and maximizing the domain classification loss. We present the generative knowledge transfer, which aligns with the first category, and introduce inherited knowledge transfer that minimizes the target domain classification loss and the discrepancy between domain feature distributions.
@@ 3 Day-to-Night Translation Datasets
In this section, we elucidate the motivation behind constructing the NightStreets and NightCities datasets. Subsequently, we delve into details regarding the composition of these datasets. It is imperative to note that these two datasets are prerequisites for our methodology.
Figure 2: Potential Datasets for Day-to-Night Translation. (a) LOL dataset [59], (b) RobotCar dataset [34], (c) night2day dataset [29], and (d) NightStreets dataset.
3.1 Motivation for Datasets Construction
@@
We present a novel observation: existing large-scale, multi-view VPR training sets, ranging from 30k [5] to an impressive 41.2M images [9], are predominantly composed of daytime images and lack nighttime representation. This omission significantly affects their recognition capabilities under nocturnal conditions, rendering them sub-optimal compared to daytime scenarios. While some day-night VPR training sets [34, 58] have demonstrated the benefits of nighttime training data for improved cross-domain recognition, they suffer from limited generalization. Two main factors contribute to this limitation: i) a consistent nighttime style across the datasets and ii) a considerably smaller scale compared to daytime datasets. We also recognize two major challenges: i) Constructing a nighttime VPR dataset that aligns with the GPS tags or 6 DoF poses of existing VPR datasets is exceptionally difficult. ii) Building a Day-Night VPR dataset from scratch, which matches the scale of current datasets, is also a daunting task. Given these challenges, we propose an indirect approach: utilizing I2I translation techniques to model daytime images in a nighttime style. Consequently, there is an urgent need for datasets encompassing both day and night scenarios for training an I2I translation network.
@@
Regrettably, our search for a suitable day-to-night I2I translation dataset was unsuccessful. As illustrated in Fig. 2 (a), most low-light enhancement datasets are primarily concerned with low-light conditions [59, 69], such as those induced by backlighting or weak exposure, rather than authentically capturing the diverse nuances of real nighttime scenes. Fig. 2 (b) showcases examples from the aforementioned day-night VPR dataset [34]. Models trained on these datasets have limited generalization capabilities for night-to-day VPR, suggesting that theoretically generated day-to-night scene data might also be inadequate. Furthermore, we observed recurring limitations such as static camera orientations due to vehicular mounting, leading to image blur. Fig. 2 (c) references several time-lapse photography datasets proposed in the previous I2I transfer research [29, 27]. However, these datasets tend to focus on distant views and skylines, deviating significantly from urban scenes, and a day-to-night time-lapse video contains higher redundancy than a few photos.
@@
Figure 3: Examples of the NightCities dataset. Only nighttime images are displayed here. From left to right, top to bottom, they represent Tokyo, Paris, Zurich, Hong Kong, Washington, San Francisco, Sydney, Munich, and Vancouver.
3.2 The NightCities Dataset
@@
Consequently, we collected a large-scale unpaired day-night urban scene image dataset named NightCities, as illustrated in Fig. 3. It encompasses 5,000 images from 60 cities globally, encapsulating diverse lighting styles, cultural vibes, and architectural features. Notably, the quantity of images aligns with representative datasets in the I2I translation domain [13, 28, 70]. The only annotation required was the categorization into day and night. We trained an unpaired I2I translation network on the NightCities dataset and used it to process a subset of the SF-XL dataset, resulting in SF-XL-NC. It consists of a training set of 59,650 night-style images as well as a validation set of 7,993 night-style queries and 8,015 database images.
3.3 The NightStreets Dataset
@@
We also built a test set for evaluating I2I translation by reorganizing query images from Tokyo 24/7 v3 [52] and Aachen Day/Night v1.1 [46], as illustrated in Fig. 2 (d). Tokyo 24/7 v3[52] comprises 375 daytime and nighttime images each, while Aachen Day/Night v1.1 [46] contains 234 daytime and 196 nighttime images. To distinguish it from the previous data sets, we call it NightStreets.

### 21 Mar 2024 v2

# NocPlace: Nocturnal Visual Place Recognition via Generative and Inherited Knowledge Transfer

@@ Abstract
Visual Place Recognition (VPR) is crucial in computer vision, aiming to retrieve database images similar to a query image from an extensive collection of known images. However, like many vision tasks, VPR always degrades at night due to the scarcity of nighttime images. Moreover, VPR needs to address the cross-domain problem of night-to-day rather than just the issue of a single nighttime domain. In response to these issues, we present NocPlace, which leverages generative and inherited knowledge transfer to embed resilience against dazzling lights and extreme darkness in the global descriptor. First, we establish a day-night urban scene dataset called NightCities, capturing diverse lighting variations and dark scenarios across 60 cities globally. Then, an image generation network is trained on this dataset and processes a large-scale VPR dataset, obtaining its nighttime version. Finally, VPR models are fine-tuned using descriptors inherited from themselves and night-style images, which builds explicit cross-domain contrastive relationships. Comprehensive experiments on various datasets demonstrate our contributions and the superiority of NocPlace. Without adding any real-time computing resources, NocPlace improves the performance of Eigenplaces by 7.6% on Tokyo 24/7 Night and 16.8% on SVOX Night.
Keywords:
Visual place recognition Domain adaption Transfer learning
@@ 1 Introduction
Figure 1: A demo for NocPlace. The center of this figure presents a map of San Francisco. 2.8 million daytime database images and 466 nighttime query images from SF-XL [8] are projected onto this map via GPS tags, corresponding to the dense cyan-blue and sparse red dots, respectively. Despite the dazzling lights and extreme darkness, our method achieves higher recall@N than previous methods.
Visual Place Recognition (VPR) is a foundational task in computer vision [5, 28, 59, 67, 68, 83] and intelligent robotics [23, 41, 43, 45, 51, 74, 42, 77, 61, 35] and is mainly used in Augmented Reality (AR) [55, 17] autonomous systems [43, 53]. VPR aims to retrieve similar database images to a given query image by comparing it against a vast set of known database images. Existing literature identifies several challenges for VPR, including database scale [8], viewpoint shifts [9, 11], repeated structures [68], structural modifications [27], occlusions [5], visual scale differences [20], illumination changes [4, 13, 43, 41, 57, 66, 67, 74, 83, 60, 39, 37, 46, 47], and seasonal transitions [66, 45, 65]. Recent VPR approaches primarily leverage large-scale datasets [8, 44, 49, 67, 76, 3] and utilize a neural network to map images into an embedding space, thereby effectively distinguishing images from different locations.
However, night-to-day VPR is constrained by the current training datasets. On the one hand, recent studies [5, 2, 11, 1, 8] leveraging Google Street View (GSV) [^1] https://developers.google.com/streetview/ have produced large-scale VPR datasets covering nearly all challenges except those in nighttime conditions. On the other hand, while some datasets specifically targeting night-to-day VPR datasets [43, 45] have been proposed and shown the benefit of incorporating nighttime data for cross-domain recognition [47, 46], their scale or data acquisition methods impede the training of a universally night-to-day VPR model. What is evident is that building nighttime datasets that align with commercial products like GSV [3] is challenging.
Due to the above restrictions, image enhancement algorithms (IE) [86, 85, 26, 25, 39] and night-to-day Image-to-Image (I2I) translation [4] may be a potential approach. These methods, however, present several limitations: i) Introducing an auxiliary model increases computational demands. ii) Most low-light enhancement models rely on images with diverse exposure rather than real night-to-day transitions. Some images do not even come from the outdoors [75]. iii) Imperfect night-to-day translations might adversely affect VPR performance [4].
In this paper, we reverse the night-to-day thinking [4] and study how to transfer night knowledge directly into neural networks. Further, we study the use of daytime knowledge to guide the training of nighttime models. As shown in Fig. 1, our method can improve the performance of nighttime VPR without consuming additional real-time computing resources.
@@
  • We propose NocPlace, a scalable Nocturnal Visual Place Recognition method that significantly narrows the VPR performance gap between daytime and nighttime domains.
  • We propose NightCities, an unpaired day-night urban scene image dataset. We utilize this dataset and an I2I translation network to transfer night knowledge to a large-scale daytime VPR dataset.
  • We propose the inherited knowledge transfer, which employs descriptors inherited from previous VPR models to guide the training of new VPR models.
  • We propose a partial-divide-and-conquer-based retrieval strategy for practical applications, which balances storage and performance.
  • Extensive experiments demonstrate that NocPlace substantially outperforms previous methods. Specifically, our VGG16-based model excels over state-of-the-art ResNet50-based models on certain test sets.
  • The source code and trained models for NocPlace are available at https://github.com/BinuxLiu/NocPlace.
@@ 2 Related Work
Visual Place Recognition can be categorized along three pivotal axes: model architecture, loss function, and dataset. About model architecture, Convolutional Neural Networks (CNNs) [5, 62] and Visual Transformers (ViTs) [69, 72, 10] emerge as the primary choices for feature extraction. These are often followed by pooling or aggregation mechanisms like NetVLAD [5] and GeM [54]. In loss functions, triplet and classification loss are frequently employed to optimize Euclidean margins for enhanced feature embeddings [10, 5]. Ge et al. [22] utilized triplet loss and network output scores as self-supervised indicators for iterative training. However, triplet loss encounters challenges with hard negative sample mining. In response, Berton et al. [8, 11] incorporated Large Margin Cosine (LMC) Loss [71] in VPR, showcasing its superiority over triplet loss. On the dataset front, we observe two primary classifications: self-built datasets and web-sourced ones. The former, such as [23, 43, 45], often exhibit small scale, fixed frontal camera view, and little nighttime data. The latter can be sub-categorized into datasets like Google Landmark [54], derived from Google Images, and those like SF-XL [8] and Pittsburgh [5], sourced from GSV [5, 1, 3, 8]. It is worth noting that the latter contains almost no night images.
@@
Night Computer Vision encompasses downstream tasks using nighttime images, which are divided into one-stage and two-stage approaches. The one-stage methods utilize nighttime data for training [15, 83]. Cui et al. [15] developed a technique for reverse ISP and dark processing, leading to an end-to-end model for object detection in dark conditions. Additionally, Xu et al. [79] presented a GAN-centric strategy for semantic segmentation using transformed nighttime imagery. Attila et al. [37] introduced the color invariant layer, reducing the day-night distribution shift in feature map activations. In contrast, two-stage approaches [4, 79] first transform nighttime images into daytime-style images before executing the downstream tasks. Anoosheh et al. [4] introduced a CycleGAN-based method to convert nighttime images for retrieval-based localization. Neubert and Schubert [47] proposed a simple algorithm to learn domain-specific descriptors from a small database. However, the two-stage methods are criticized for their limited generalizability [4] and heightened computational demands. Meanwhile, the one-stage approach faces challenges due to the lack of extensive training datasets. In response, we propose a pioneering framework by converting daytime images to nighttime versions for enhanced VPR training.
@@
Image-to-Image Translation pertains to the intersection of vision and graphics, aiming to establish a mapping between an input and an output image [89]. These methods are primarily categorized based on the nature of their training sets: pixel-level matched (paired) I2I [31] or unpaired I2I [89, 81]. Collecting extensive paired day-night images in real-world scenarios is notably challenging. Hence, our focus is on the latter training set type. Two predominant strategies for unpaired I2I translation exist bi-directional mapping and uni-directional mapping. Bi-directional mapping, encompassing CycleGAN [89] and DualGAN [81], is based on the cycle-consistency constraint, ensuring the transposed image can undergo reverse mapping for reconstruction. The bi-directional mapping can be overly restrictive, prompting the evolution of uni-directional mapping methods. For instance, DistanceGAN [19] conserves inter-class distances within domains, while GCGAN [7] upholds geometry consistency between inputs and outputs. The CUT [50] amplifies mutual information between inputs and outputs via contrastive loss. The diffusion model [88] recently has shown exceptional performance. However, we found that it consumes considerable computational resources and is unsuitable for generating large-scale data. After thoroughly comparing different methods, we use NEG-CUT [73] for nighttime data generation.
@@
Knowledge Transfer can be categorized in various ways based on the problems it addresses, such as domain adaptation [84], multi-task learning [87], meta-learning [70], and more. This paper focuses on domain adaptation, aiming to adapt models trained in one domain (source) to perform optimally in another domain (target) where the data distribution might differ. Previous Night-to-Day VPR studies primarily focused on labeled target domains, a relatively straightforward problem [77, 80, 52]. Cao et al. [52] incorporated Canny edge descriptors and the Wasserstein distance measure for domain adaptation in VPR tasks. When dealing with labeled source and unlabeled target domains, prior research can be broadly divided into two categories: i) generation of target domain data, as exemplified by SimGAN [63], which transfers samples from the source to the target domain while preserving consistent labels. ii) adversarial learning of domain-invariant features. Ganin et al. introduced DANN [21], which learns domain-invariant feature distributions by minimizing the source domain classification loss and maximizing the domain classification loss. We present the generative knowledge transfer, which aligns with the first category, and introduce inherited knowledge transfer that minimizes the discrepancy between domain feature distributions.
Figure 2: Potential Datasets for Day-to-Night Translation. From left to right, they represent (a) LOL [75], (b) RobotCar [43], (c) day2night [36], and (d) Tokyo 24/7 [67].
@@ 3 Day-to-Night Translation Datasets
This section introduces the motivation behind constructing the NightStreets and NightCities datasets. Subsequently, we delve into details regarding the composition of these datasets. It is important to note that these datasets are prerequisites for our methodology and evaluation.
@@
Motivation for Datasets Construction. We present an observation: existing large-scale, multi-view VPR training sets, ranging from 30k images [5] to a remarkable 41.2M images [8], mainly consist of daytime images and lack nighttime images. This omission significantly affects the representation capabilities of previous work under nocturnal conditions, making them less optimal than in daytime scenarios. While some day-night VPR training sets [43, 74] have demonstrated the benefits of nighttime training data for improved cross-domain recognition, they suffer from limited generalization. Two main factors contribute to this limitation: i) the presence of only a single nighttime style in some datasets [43] and ii) a considerably smaller scale than daytime datasets. We also recognize two significant challenges: i) Constructing a nighttime VPR dataset that aligns with the GPS tags or 6 DoF poses of existing VPR datasets is exceptionally difficult. ii) Building a Day-Night VPR dataset from scratch, which matches the scale of current datasets, is also a daunting task. Given these challenges, we propose an indirect approach: utilizing I2I translation techniques to convert daytime images into a nighttime style. Consequently, there is an urgent need for datasets encompassing both day and night scenarios for training an I2I translation network.
@@
Regrettably, our search for a suitable day-to-night I2I translation dataset was unsuccessful. As illustrated in Fig. 2 (a), most low-light enhancement datasets are primarily concerned with low-light conditions [75, 86], such as those induced by backlighting or weak exposure, rather than capturing the diverse nuances of real nighttime scenes. Fig. 2 (b) showcases examples from the aforementioned day-night VPR dataset [43]. Models trained on these datasets have limited generalization capabilities for night-to-day VPR, suggesting that theoretically generated day-to-night scene data might also be inadequate. Furthermore, we observe recurring limitations in autonomous driving datasets, such as image blur and single camera orientation. Fig. 2 (c) references several time-lapse photography datasets proposed in the previous I2I transfer research [36, 31]. However, these datasets tend to focus on distant views and skylines, deviating significantly from urban scenes, and a day-to-night time-lapse video contains higher redundancy than a few photos. As shown in Fig. 2 (d), the night test sets of VPR can be used to train I2I, but it lacks fairness because it may lead to implicit data leakage.
@@
Figure 3: Examples of the NightCities dataset. Only nighttime images are displayed here, and daytime images are shown in the supplementary material. The images are arranged from left to right and top to bottom, representing the following cities in order: Tokyo, Hong Kong, San Francisco, Paris, Milan, Sydney, Munich, Zurich, and Vancouver.
@@
The NightCities Dataset. Consequently, we collected a large-scale unpaired day-night urban scene dataset named NightCities, as illustrated in Fig. 3. It encompasses 5,000 images from 60 cities globally, encapsulating diverse lighting styles, cultural vibes, and architectural features. Notably, the quantity of images aligns with representative datasets in the I2I translation domain [14, 34, 88]. The only annotation required was the categorization into day and night. We trained an unpaired I2I translation network on NightCities and used this network to process a subset of SF-XL, resulting in SF-XL-NC. It consists of a training set of 59,650 night-style images, a validation set of 7,993 night-style queries, and 8,015 database images.
@@
The NightStreets Dataset. We also built a test set for evaluating I2I translation by reorganizing query images from Tokyo 24/7 v3 [67] and Aachen Day/Night v1.1 [58], as illustrated in Fig. 2 (d). Tokyo 24/7 v3 [67] comprises 375 daytime and nighttime images each, while Aachen Day/Night v1.1 [58] contains 234 daytime and 196 nighttime images. To distinguish it from the NightCities dataset, we call it NightStreets.